### v1

# Towards Interpretable Deep Generative Models via Causal Representation Learning

Abstract
Recent developments in generative artificial intelligence (AI) rely on machine learning techniques such as deep learning and generative modeling to achieve state-of-the-art performance across wide-ranging domains. These methods’ surprising performance is due in part to their ability to learn implicit “representations” of complex, multi-modal data. Unfortunately, deep neural networks are notoriously black boxes that obscure these representations, making them difficult to interpret or analyze. To resolve these difficulties, one approach is to build new interpretable neural network models from the ground up. This is the goal of the emerging field of causal representation learning (CRL) that uses causality as a vector for building flexible, interpretable, and transferable generative AI. CRL can be seen as a culmination of three intrinsically statistical problems: (i) latent variable models such as factor analysis; (ii) causal graphical models with latent variables; and (iii) nonparametric statistics and deep learning. This paper reviews recent progress in CRL from a statistical perspective, focusing on connections to classical models and statistical and causal identifiablity results. This review also highlights key application areas, implementation strategies, and open statistical questions in CRL.
Keywords: generative models, causality, machine learning, deep learning, latent variable models
1 Introduction
Generative AI has achieved state-of-the-art performance on a variety of tasks across numerous domains; for example, predicting protein folding (Jumper et al.,, 2021), generating new candidate molecules for drug discovery (Merchant et al.,, 2023), text-to-image generation (Betker et al.,, 2023) and test taking (OpenAI,, 2024). This performance has been driven by machine learning (ML) techniques such as deep learning and generative modeling.
These methods’ surprising performance is due in part to their ability to learn implicit “representations” of complex, multi-modal data. A representation is simply a latent numerical summary of data that encodes semantically meaningful features, and can be interpreted analogously to latent factors in factor analysis. Indeed, this interpretation is explicit in early work on probabilistic machine learning (Roweis and Ghahramani,, 1999; Lawrence,, 2005; Goodfellow et al.,, 2012). This early work focused on latent variable models such as Gaussian process latent variable models (GPLVMs), mixtures, and hierarchical Bayesian models, however, the past decade has witnessed a shift towards deep learning, where nonlinearities are captured with deep neural networks of essentially arbitrary complexity (Rezende et al.,, 2014; Kingma and Welling,, 2014; Ranganath et al.,, 2014).
Unfortunately, as is well-known, neural networks are “black boxes” that obscure such data representations, making them difficult to even define concretely, let alone interpret, visualize or analyze. Without understanding how ML algorithms are learning and manipulating data representations, it is difficult to trust that these algorithms are performing as they should. Moreover, understanding such algorithms is crucial to gaining scientific insight into the data problem at hand; performance in prediction and generation is not enough.
To resolve these difficulties in understanding ML algorithms, there are two main directions in which to proceed. The first direction is to develop tools to interpret and analyze pre-trained models (often called “post-hoc explainability”). Such post-hoc tools and analyses have already provided insight into image and language models (Selvaraju et al.,, 2017; Templeton et al.,, 2024). However, a problem with such post-hoc explainability methods is that they may not accurately reflect what the original model computes (Rudin,, 2019); moreover, many explainability methods are provably no better than random guessing at inferring model behavior (Bilodeau et al.,, 2024). As an alternative, the second direction is to build new models which have more transparent and interpretable internal mechanisms. By building interpretable models from “the ground up”, the hope is to obtain more confidence in and understanding of their conclusions.
[…]
(a) Example 1: Gene expression data. Left: Graphical model where the expression of gene $x_{j}$ is drawn from latent low-dimensional factors $z_{k}$ with dependence among the latent $z_{k}$. Right: Matrix of gene expression data from Zeisel et al., (2015) (subset of 500 cells by 300 genes).
[…]
(b) Example 2: Cell image data. A generative model for images of perturbed cells. In each image, different compounds have been added to the cells, represented as a perturbation of the latent features of the cell. Images from Bray et al., (2017) (top: CIL45758; bottom: CIL45766).
Figure 1: Examples of generative models. The causal representation learning problem is to recover the latent factors along with the dependence structure, both of which are assumed unknown, from the observed features. Although the latent factors are low-dimensional, the nonlinear mapping $f$ can be estimated using a deep neural network with many more hidden units than the number of latent factors.
One first step to building more interpretable models is to consider a generative model where the latent representations are explicitly parameterized, as in Figure 1 (a more formal development will be deferred until later sections). Here, the training data is given by a high-dimensional random vector $\bm{x}_{i}$, and we seek a latent “representation” of $\bm{x}_{i}$ via some hidden variables $\bm{z}_{i}$. We assume that $\bm{x}_{i}$ is generated from $\bm{z}_{i}$ via an (unknown) function $f$. In this setting, $\bm{z}_{i}$ encodes the concept of a “learned representation” mathematically. As an example, $\bm{x}_{i}$ may be a vector of gene expression values, and the $\bm{z}_{i}$ may represent different biological processes. In this generative model, switching “on” a dimension of $\bm{z}_{i}$ will then impact the expression levels of genes associated with that biological process. As another example, $\bm{x}_{i}$ may be pixels of images and $\bm{z}_{i}$ may correspond to image features such as content and style. In an ideal interpretable model, changing the dimension of $\bm{z}_{i}$ corresponding to “style” will generate images with different stylistic choices while keeping the content of the image fixed.
Many modern deep generative models are explicitly of this form, including autoencoders, diffusions and flows. The study of such generative models, however, has focused mainly on the problem of generation; equivalently, sampling or density estimation. That is, the goal is simply to produce realistic samples from complex distributions over text, images, audio, and/or video. Crucially, as is well-known, it is possible to sample from a distribution as a black-box without any conceptual understanding of underlying mechanisms or processes. As a result, the problem of understanding generative models, especially in terms of estimation of the parameters $(\bm{z}_{i},f)$, has been less well studied. This distinction between density and parameter estimation is not unique to generative models, and parallels the “two cultures” of statistical modeling, namely prediction-focused and inference-focused modeling (Breiman,, 2001); see Section 5 for more discussion.
To understand and interpret generative models as in Figure 1, parameter estimation is crucial. A key technical challenge in this endeavor is that generative models with different parameters can still produce identical density estimates, despite their internal mechanisms being completely different. Phrased more formally, the parametrization of generative models given in Figure 1 is non-identifiable. Moreover, this non-identifability makes it difficult to certify the behavior of a generative model using post-hoc explainability techniques. To render models identifiable, we need to either a) restrict the form of the model, or b) gain access to richer data modalities arising from different environments.
In classical analogues of representation learning, such as factor models and independent component analysis, identifiability conditions and consequently interpretable parameter estimation are well-studied. (Later, in Section 3, we outline the connections between generative models and classical linear factor models in particular.) However, such classical approaches with restrictive parametric models are often insufficient to model complex unstructured and multimodal datasets that proliferate in modern applications of AI. In these settings, more flexible models that avoid these restrictions are needed. Thus the focus shifts from model restrictions to richer data modalities. Further, there is increased demand for models to transfer well to a variety of downstream tasks (as opposed to specializing to specific, well-defined tasks a priori, which is often the case for classical models).
Motivated by these goals of flexibility, interpretability, and transferability, the field of causal representation learning (CRL) has emerged to tackle these issues by bringing the flexibility of deep learning together with the interpretability and transferability of causality. It is widely believed that modeling and capturing causal relationships results in stable, robust, and invariant ML models that are easier to interpret. Further, the value of causal models for stable generalization and transfer to unseen tasks is widely recognized in the ML community.
2 Motivating examples
The idea of identifying latent factors with causal implications is at the heart of scientific discovery. To help ground our discussion, the following two examples serve to motivate some of the key issues behind CRL.
Motivating example 1: Genomics
Perturb-Seq (Dixit et al.,, 2016) combines single-cell RNA sequencing with CRISPR-based perturbations of specific genes. The data consists of an observational dataset $\bm{X}\in\mathbb{R}^{N\times D}$ of $N$ cells and $D$ genes, and interventional datasets $\bm{X}^{(e)}\in\mathbb{R}^{N_{e}\times D}$, $e\in\mathcal{E}$. In each intervention $e$, a particular gene is targeted by CRISPR perturbation. Perturbation types include gene knockouts (deletion or disruption of target genes via CRISPR-Cas 9, Cong et al.,, 2013), gene knockdowns (reduction in gene expression via CRISPRi, Qi et al.,, 2013) or gene activation (CRISPRa to enhance gene expression, Gilbert et al.,, 2014). See Figure 1(a).
These interventional data provide an opportunity to better understand gene regulatory networks. One approach to analyzing this data is to analyze the gene expression levels directly, using linear regression on the perturbation indicator or causal discovery algorithms. However, such approaches do not take advantage of the latent low-dimensional structure in the data. A potentially more efficient approach to causal discovery is to reduce the dimension of the data and consider the causal graph of the latent space. For example, if a set of genes exhibit coordinated expression, representing a gene program, we can consider these genes as a single causal variable and seek to interpret its biological function. See Zhang et al., (2023) for a real data example with Perturb-Seq data in the context of CRL.
Motivating example 2: Unstructured data
It is becoming increasingly common to analyze unstructured data in the form of images, text, and audio. A key feature of this data is that the measurements—pixels in images, tokens in text, and samples in audio—are typically transformed before they are analyzed. In images, for example, interpreting raw pixels is not useful and individual pixels do not carry causal content (e.g. intervening on a pixel is generally not meaningful in applications). This is also becoming increasingly common in biological applications with images (e.g. super-resolution microscopy, cryo-EM, and cell painting) and text (e.g. sequencing data). See Figure 1(b).
More precisely, consider image data $\bm{X}\in\mathbb{R}^{N\times D}$ of $N$ images and $D$ pixels, and interventional datasets $\bm{X}^{(e)}\in\mathbb{R}^{N_{e}\times D}$, $e\in\mathcal{E}$. In biological applications, each intervention $e$ may correspond to CRISPR knockouts (Jain et al.,, 2024) or drug administration (Chandrasekaran et al.,, 2021). More broadly, in general imaging and vision applications, interventions $e$ can correspond to different forms of data augmentation, e.g. images can be rotated, translated, rescaled, colour corrected, or texturized. In these applications, the effects of data augmentation or a gene knockout cannot be detected at the individual pixel level, and instead these effects act on latent factors that must be inferred from the data. See Von Kügelgen et al., (2021); Brehmer et al., (2022) for real data examples with 3D images in the context of CRL.
In both of these examples, learning latent causal factors can aid in interpretability, computational efficiency, and potentially improve power to detect effects via data pooling, but there are a number of challenges. Are these latent abstractions identifiable and interpretable? When is it possible to obtain the causal graph of the latent variables? These are the goals of causal representation learning.
3 Linear Representation Learning
[…]
$\bm{x}_{i}$ $=\bm{B}\bm{z}_{i}+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\sim N(\bm{0},\bm{\Sigma}),\quad\bm{z}_{i}\sim p_{\bm{z}}(\bm{z}_{i}),$ (3.1)
where $\bm{B}\in\mathbb{R}^{D\times K}$ is the factor loadings matrix and $\bm{\Sigma}=\text{diag}(\sigma_{1}^{2},\dots,\sigma_{D}^{2})$. This is a special case of the general model in Figure 1 where the transformation $f$ is linear.
Figure 2: A simulated dataset from the linear factor analysis model with $N=100$ samples, $D=1956$ features, and $K=5$ factors. (a) The data $\bm{X}$, latent factors $\bm{Z}$ and loadings $\bm{B}$; (b) The empirical covariance matrix $\bm{X}^{\top}\bm{X}$.
A common goal of linear factor analysis is to find “scientifically meaningful” groups of features. In the linear factor analysis model, “scientifically meaningful” is operationalized as a group of features which covary. These covarying groups can be inferred from the loadings matrix, $\bm{B}$. Specifically, the $k$th group of features corresponds to the non-zero elements of $\bm{B}_{\cdot,k}$, the $k$th column of the loadings matrix (Figure 2(a)).
Identifiability.  The goal of finding meaningful factors is hindered by the well-known issue of identifiability in (3.1). Specifically, the linear factor analysis model is not identifiable when $p_{\bm{z}}\sim N(\bm{0},\bm{I})$ To see this, suppose $\bm{z}_{i}$ and $\bm{B}$ are the true parameters that generate $\bm{x}_{i}$. Then, for any rotation matrix $\bm{P}$ we have:
$\bm{x}_{i}=\bm{B}\bm{z}_{i}+\bm{\varepsilon}_{i}=\bm{B}\bm{P}\bm{P}^{\top}\bm{z}_{i}+\bm{\varepsilon}_{i}=\widetilde{\bm{B}}\widetilde{\bm{z}}_{i}+\bm{\varepsilon}_{i}.$
The above reparameterization is valid as the standard Gaussian is invariant under rotation: $\bm{P}^{\top}\bm{z}_{i}\sim N(0,I)$. Consequently, the model is not identifiable. This lack of identifiability is problematic for interpreting groups of features via the loadings matrix $\bm{B}$: There are an infinite number of loadings matrices with the same likelihood; based on the data alone, we cannot determine the true loadings matrix.
Figure 3: Linear factor analysis is not identifiable; applying any rotation matrix $\bm{P}$ leads to another model with equivalent likelihood.
The identifiability problem can be resolved by placing further assumptions on the data generating process. Typically, further assumptions are placed on either (i) the class of loadings matrices, (ii) the distribution $p_{\bm{z}}$, or both. The goal is to specify assumptions on $\bm{z}_{i}$ and $\bm{B}$ that are not preserved under rotation. Then, $(\bm{P}^{\top}\bm{z}_{i},\bm{B}\bm{P})$ will not satisfy these assumptions for any nontrivial rotation $\bm{P}$, and thus will not be equivalent solutions.
3.1 Identifiability assumptions on the loadings matrix
Triangular loadings matrices.  One early approach to identify the linear factor analysis model was to constrain the class of loadings matrices to be lower triangular matrices with positive diagonal elements (PLT) (Anderson and Rubin,, 1956). There are no rotation matrices that preserve this PLT structure except the identity, so $\bm{B}$ is identifiable. The PLT assumption is commonly used in the statistical factor analysis literature (Carvalho et al.,, 2008; Frühwirth-Schnatter and Lopes,, 2010).
While the PLT identifies $\bm{B}$, it does so at the expense of “fixing” the factors to the first $K$ dimensions of $\bm{X}$. That is, with the triangular assumption, the generative model is:
$x_{i1}$ $=b_{11}z_{i1}+\varepsilon_{i1},$
$x_{i2}$ $=b_{21}z_{i1}+b_{22}z_{i2}+\varepsilon_{i2}$
$\vdots$
$x_{iK}$ $=b_{K1}z_{i1}+\cdots b_{KK}z_{iK}+\varepsilon_{iK},\quad i=1,\dots,N.$
The first factor dimension $\bm{z}_{\cdot 1}$ then corresponds to the first observed feature $\bm{x}_{\cdot 1}$; the second factor dimension corresponds to a linear combination of the first two observed features, and so on. This correspondence may be problematic in applications where the first $K$ observed features do not have special status for defining the factors.
Sparsity and anchor features.
More recently, Bing et al., (2020) proposed a method which learns which observed features should be used as “anchors” for the factor dimensions. Bing et al., (2020) prove their method identifies the loadings matrix with the key assumption of “pure variables”, or what we call “anchor features”. An anchor feature for a particular factor is an observed feature that depends only on this factor. In other words, no other factor influences this feature. The anchor features assumption of Bing et al., (2020) then demands that for every factor dimension, there are at least two anchor features. The anchor features do not need to be known in advance; these features have observable implications in the data covariance and thus can be inferred (see Figure 2(b) for a visualization).
The anchor features assumption means that the true loadings matrix $\bm{B}$ contains two submatrices that are the identity (after row and column permutation):
$\bm{B}^{\top}=\begin{pmatrix}\bm{I}_{K\times K}&\bm{I}_{K\times K}&\widetilde{\bm{B}}^{\top}\end{pmatrix}.$ (3.2)
This structure on $\bm{B}$ again eliminates all non-trivial rotations because this particular sparsity pattern is not invariant to rotation.
The anchor feature assumption has also been utilized for identification in related latent variable models. For example, this is analogous to the “anchor word” assumption in topic models (Arora et al.,, 2013). This is also related to the separability condition (Donoho and Stodden,, 2003) in the non-negative matrix factorization (NMF) literature.
The anchor feature and separability assumptions also help explain the empirical success of sparse Bayesian factor analysis. In Bayesian factor analysis, sparsity-inducing priors are often placed on the loadings matrix $\bm{B}$; these sparse models often yield interpretable groups of features in real data settings (Knowles and Ghahramani,, 2011; Bhattacharya and Dunson,, 2011; Ročková and George,, 2016; Wang and Stephens,, 2021). These sparsity-inducing priors will place higher probability on loadings matrices that satisfy the anchor feature assumption assumption than dense loadings matrices.
3.2 Identifiability assumptions on the distribution of the factors
Heavy-tailed factors.  Rohe and Zeng, (2023) show that the factors are identifiable if their distribution belongs to a class of heavy-tailed distributions. These heavy-tailed distributions are not rotationally invariant, thus limiting the set of equivalent-likelihood solutions. [^1] Maxwell’s theorem states that the multivariate Gaussian is the only rotationally invariant distribution of independent variables (Maxwell,, 1860 and III, 4 in Feller,, 1971.) The heavy-tailed assumption on the factors also has precedent in the literature on independent component analysis (ICA, Comon,, 1994). Another key aspect of Rohe and Zeng, (2023) is that they provide theoretical justification for a long-used technique: principal component analysis (PCA) followed by a Varimax rotation (Kaiser,, 1958). Specifically, Rohe and Zeng, (2023) prove this “vintage” procedure consistently estimates the latent factors.
Correlated factors.
Typically the latent factors are assumed to be uncorrelated, but recent work has begun to investigate the case where factors are correlated: Bing et al., (2020) allow for correlated factors, as long as the variance of each factor is greater than its covariance with any other factor; Rohe and Zeng, (2023) implicitly allow for correlated factors by estimating a “middle matrix” between the loading and the factor matrices.
3.3 Beyond classical factor analysis
In the machine learning literature, similar identifiability conditions have been proposed using the language of graphical models and ICA. For example:
  • The PLT assumption is equivalent to assuming a known causal ordering on the factors;
  • Anchor features are equivalent to assuming a particular sparsity pattern on the graphical model implied by the loadings matrix $\bm{B}$;
  • The use of heavy-tailed factors is closely related to non-Gaussianity assumptions that have been used to identify ICA models;
  • Correlated factors arise naturally with causally related latent factors, and assumptions on correlated factors are closely related to causal restrictions on the latent factors.
Silva et al., (2006) recognized this connection with causal models, and presented a framework for learning causal structure in linear latent variable models. For example, can we determine which biological processes are causally related? In essence, this is the goal of causal representation learning: How do we learn latent factors (i.e. representations), and learn the causal relationships between them? To study such questions, we need some concepts some the graphical approach to causality (Pearl,, 2000), which we now review.
4 Causal Models and Latent Variables
Evidently, there is a close connection between classical identifiability assumptions in factor analysis and graphical assumptions that arise in machine learning and causal applications. In fact, the usual factor model is equivalent to a particular type of graphical model, sometimes called a measurement model (Silva et al.,, 2006), where there are no edges between observed variables, or from observed variables to latent variables. This relates to the factor analysis model where the $\bm{x}_{i}$ are independent conditional on $\bm{z}_{i}$.
In order to keep this review self-contained, in this section we review some minimal background on graphical models, with a focus on the linear setting to allow for easy comparison with the previous section. Readers already familiar with graphical models will immediately recognize that everything extends to nonlinear models and in subsequent sections, we will consider general nonlinear models.
4.1 Graphical approach to causality
The graphical approach to causality is built upon the fundamental concept of a Bayesian network. A Bayesian network (BN) over a random vector $\bm{x}_{i}=(\bm{x}_{i1},\ldots,\bm{x}_{iD})\in\mathbb{R}^{D}$ with joint distribution $P$ is any directed acyclic graph (DAG) $\mathsf{B}=(\mathsf{V},\mathsf{E})$ with $\mathsf{V}=\bm{x}_{i}$ such that $P$ factorizes over $\mathsf{B}$, i.e. it satisfies the Markov condition or is Markov to $\mathsf{B}$:
$P(\bm{x}_{i})=\prod_{j=1}^{D}P(\bm{x}_{ij}\,|\,\operatorname{pa}_{\mathsf{B}}(j)).$ (4.1)
[…]
We are interested in models that contain both observed and hidden components: $(\bm{x}_{i},\bm{z}_{i})=(\bm{x}_{i1},\dots,\bm{x}_{iD},\bm{z}_{i1},\dots,\bm{z}_{iK})$. In this case, the DAG $\mathsf{B}$ and the factorization (4.1) extends over the joint random vector $(\bm{x}_{i},\bm{z}_{i})$. Without some type of structural assumption on $\mathsf{B}$, the types of structures that can be represented by a Bayesian network with latent variables are known to be quite exotic, and have been the subject of much research (Evans,, 2016; Evans et al.,, 2018; Richardson et al.,, 2017). We will focus our review on the case of factor structures, defined below, in order to emphasize the connection with factor analysis, although one could of course study more exotic latent structures.
Factor models as in (3.1) implicitly impose structural assumptions in terms of the DAG $\mathsf{B}$ (Drton et al.,, 2007, 2023). That is, if $\mathsf{B}$ is a BN for a factor model over $(\bm{x}_{i},\bm{z}_{i})$, then $\mathsf{B}$ has a factor structure defined as follows:
Definition 1.
A DAG $\mathsf{B}$ is said to have a factor structure if there are no edges between observed variables or from observed variables to latent variables. A distribution over $(\bm{x}_{i},\bm{z}_{i})$ is said to follow a factor model if $(\bm{x}_{i},\bm{z}_{i})$ is Markov to a DAG $\mathsf{B}$ with a factor structure.
This definition amounts to the assumption that $\bm{x}_{i}$ are conditionally independent given $\bm{z}_{i}$, as in the factor analysis model. In fact, if we assume linear relationships, Definition 1 is equivalent to (3.1). Figure 1 depicts two prototypical examples of a factor model.
4.2 Latent causal graphs
$z_{1}$$z_{2}$$z_{3}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$x_{7}$$\mathsf{G}$
Figure 4: Decomposition of the full causal factor model $\mathsf{B}$ over $(\bm{x}_{i},\bm{z}_{i})$ into a latent causal graph $\mathsf{G}$ over $\bm{z}_{i}$ and a bipartite graph (dashed arrows) from $\bm{z}_{i}$ to $\bm{x}_{i}$.
The factor model makes explicit that we are not interested in causal relationships between observables—it is not $\bm{x}_{i}$ itself that we are interested in modeling causally, but rather the latent structure over $\bm{z}_{i}$. The primary causal relationships of interest involve only the hidden variables. Thus, all of the modeling emphasis is shifted over to the latent factors. The goal is to use $\bm{x}_{i}$ to infer a latent causal structure: This involves learning both the latent variables $\bm{z}_{i}$ and their causal relationships.
More formally, factor models always decompose into two parts (Figure 4): A bipartite component that involves only directed edges from $\bm{z}_{i}\to\bm{x}_{i}$, and a latent causal graph $\mathsf{G}$ that involves only directed edges between the latent variables $\bm{z}_{i}$. Thus, the full causal graph over $(\bm{x}_{i},\bm{z}_{i})$ jointly is denoted by $\mathsf{B}$, whereas the latent causal graph over $\bm{z}_{i}$ only is denoted by $\mathsf{G}$. We omit reference to the bipartite graph since this is not the main focus. See Figure 4.
$z_{1}$$z_{2}$$z_{3}$$z_{4}$222233-3- 333331111
[…]
Figure 5: An example of a linear structural equation model over latent variables $\bm{z}_{i}$.
In the linear case (also known as a linear structural equation model, or SEM), the model for $p_{\bm{z}}(\bm{z}_{i})$ becomes
$\bm{z}_{i}=A^{T}\bm{z}_{i}+\varepsilon,\quad A=(a_{ij})\in\mathbb{R}^{K\times K},\quad\varepsilon_{j}\perp\!\!\!\perp\operatorname{pa}_{\mathsf{G}}(j).$ (4.2)
Then we can associate to $\bm{z}_{i}$ a latent causal graph $\mathsf{G}$ by reading off the nonzero entries of the matrix $A$: Simply define $\mathsf{G}=(\mathsf{V},\mathsf{E})$ by
[…]
See Figure 5 for an example SEM. It turns out that as long as $\mathsf{G}$ is acyclic, then $\mathsf{G}$ is a Bayesian network (i.e. a DAG) for $\bm{z}_{i}$. In other words, the matrix $A$ explicitly encodes a Bayesian network for $\bm{z}_{i}$.
4.3 Interventions and causality
The edges in a DAG can be interpreted as indicating the causal dependencies in a system: The edge $\bm{z}_{ij}\to\bm{z}_{ik}$ indicates that $\bm{z}_{ik}$ depends causally on $\bm{z}_{ij}$ in the sense that changes to $\bm{z}_{ij}$ propagate to $\bm{z}_{ik}$. Returning to the linear SEM example in (4.2), such causal dependencies can be seen by writing the model component-wise:
$\bm{z}_{ik}=\sum_{j=1}^{K}a_{jk}\bm{z}_{ij}+\varepsilon_{k}.$ (4.4)
Explicitly, $\bm{z}_{ik}$ is a noisy function of its parents.
Interventions and independent mechanisms.
This causal interpretation allows us to reason about how the system changes when we externally manipulate the values of certain variables, i.e. an intervention. An intervention occurs when we fix the value of a certain variable (or a set of variables) to a particular value. When this happens, the causal parents of the intervened variable are ignored, akin to deleting any incoming edges to this node.
To make this concrete, consider again the linear SEM in Figure 5. In equilibrium (in particular, without any external interventions) the value of $\bm{z}_{i3}$ is determined by the values of $\bm{z}_{i1}$ and $\bm{z}_{i2}$. The value of $\bm{z}_{i2}$, in turn, is determined by $\bm{z}_{i1}$, whereas $\bm{z}_{i1}$ is an independent exogenous variable that is not determined by any other $\bm{z}_{ik}$. Now suppose we intervene on $\bm{z}_{i2}$: Then the edge from $\bm{z}_{i1}\to\bm{z}_{i2}$ is broken, that is, the value of $\bm{z}_{i1}$ is no longer relevant for determining $\bm{z}_{i2}$: The value of $\bm{z}_{i2}$ is fixed by external manipulation. This then propagates down to $\bm{z}_{i3}$, whose value is determined by the (unintervened) value of $\bm{z}_{i1}$ and the intervened value of $\bm{z}_{i2}$. The result is a new linear SEM
$\bm{z}_{i1}$ $=\varepsilon_{1}$
$\bm{z}_{i2}$ $=\bm{z}_{i2}^{*}$
$\bm{z}_{i3}$ $=-3\bm{z}_{i1}+3\bm{z}_{i2}^{*}+\varepsilon_{3}$
$\bm{z}_{i4}$ $=\bm{z}_{i3}+\varepsilon_{4},$
where $\bm{z}_{i2}^{*}$ is the value that $\bm{z}_{i2}$ is manipulated to. To denote the new distribution of $\bm{z}_{i3}$ given this intervention on $\bm{z}_{i2}$, the convention is to use do-notation:
$P(\bm{z}_{i3}|\bm{z}_{i1},\mathrm{do}(\bm{z}_{i2}=z_{2}^{*})).$
Implicit in this formalization of interventions is the principle of modularity or independent mechanisms: Intervening on a node changes the local conditional distribution of that node, but leaves the remaining conditional distributions unchanged. See Peters et al., (2017) (Chapter 2) for more discussion on the principle of independent mechanisms.
Latent interventions.
The factor model (Definition 1) makes explicit that intervening on the observed features $\bm{x}_{i}$ is not meaningful. Instead, we are interested in intervening on latent variables, which might seem strange at first glance. After all, we do not directly observe latent variables. In many applications, however, this is quite natural: For example, recall the motivating examples in Section 2.
When the observed features are pixels (Figure 1(b)), intervening on individual pixels is not meaningful outside of toy applications. In practice, however, we can obtain additional images or data augmentations in which the object of the image is translated or rotated, or the background is changed. Each of these changes can be viewed as an intervention. However, it is difficult to model such interventions at the pixel level; e.g. in different images, different subsets of pixel values correspond to the background. A solution to this problem is to learn a latent representation $\bm{z}_{i}$ of each image $\bm{x}_{i}$ where the dimensions $\bm{z}_{ik}$ have a consistent interpretation for all $i$; e.g. $\bm{z}_{ik}$ corresponds to the rotation of the images. This type of data is readily available in applications, and can even be easily generated with the aid of simulators (e.g. Von Kügelgen et al.,, 2021; Brehmer et al.,, 2022).
In genomics, the observations are gene expression levels (Figure 1(a)). Genes may coordinate as part of a biological pathway; collectively, these genes may induce effects in other variables. As such, it may be more efficient to aggregate these gene subsets into single causal variables representing gene programs or biological processes. Other examples of latent interventions include medical treatments (drugs, surgeries), radiation exposure, or other epigenetic factors (DNA methylation, histone modification, environmental factors).
We refer to an intervention where a single dimension of $\bm{z}$ is intervened upon as a single-node intervention. For a linear SEM, we formalize this as follows.
Definition 2 (Single-node latent interventions).
A interventional distribution $p_{\bm{z}^{(e)}}$ is a single node intervention of the latent distribution $p_{\bm{z}}$ if there is a single target node $t(e)\in[K]$, and the assigning equation of $z^{(e)}_{i,t(e)}$ is:
$z_{i,t(e)}^{(e)}$ $=(A^{(e)}\bm{z}_{i}^{(e)})_{t(e)}+\varepsilon_{i,t(e)}+c^{(e)}$ (4.5)
while leaving all other equations $\bm{z}_{i,k}$ for $k\neq t(e)$ unchanged. We assume $a_{t(e),k}^{(e)}\neq 0$ only if $k\to t_{e}$ in $\mathsf{G}$; that is, no parents are added. The variable $c^{(e)}$ denotes a shift of the mean; the variance of the noise $\varepsilon_{i,t(e)}$ may also be different to that of $p_{\bm{z}}$. An intervention is called perfect if $a_{t(e),k}^{(e)}=0$ for all $k$ i.e. all connections to former parents are removed.
Intervention targets can be either known or unknown: When intervention targets are known, there is no ambiguity about the mapping $t(e)$ defined above. A considerably more challenging setting arises when intervention targets are unknown, meaning $t(e)$ is not known. In the setting of CRL, since the intervention targets are latent and unobserved, these interventions must also be unknown out of necessity—in fact, we may not even know the correct number of latent factors to begin with.
This leads to additional complications in CRL: Although there is a growing literature on learning under unknown interventions (Eaton and Murphy,, 2007; Kocaoglu et al.,, 2019; Jaber et al.,, 2020; Squires et al.,, 2020; Castelletti and Peluso,, 2022; Hägele et al.,, 2022; Faria et al.,, 2022), this line of work focuses on interventions on measured variables (i.e. $\bm{x}_{ij}$). When interventions are both unknown and latent (i.e. $\bm{z}_{ij}$), much less is known, and this is a developing area of research driven by problems in CRL; e.g. Jiang and Aragam, (2023).
5 What is causal representation learning?
With these preliminaries on factor models and causality out of the way, we are ready to discuss CRL in more precise terms.
Simply put, CRL takes a modern view on factor analysis, driven by the success of generative AI and guided by three core principles as outlined in the introduction: Flexibility, interpretability, and transferability.
  • Flexibility is achieved by allowing (a) correlations between latents, and (b) allowing for general nonlinear relationships through arbitrarily complex neural networks and deep learning;
  • Interpretability is achieved by enforcing sparsity (to simplify the model) and causality (to allow for an interventionist interpretation of the latent factors) through the use of causal graphical models;
  • Transferability is achieved through the use of causal models, which are widely believed to yield stable, invariant, transferable predictions across changing contexts and environments.
To capture these three core principles, CRL begins with a general nonlinear factor model
$\bm{x}_{i}=f(\bm{z}_{i})+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}p(\bm{\varepsilon}),\quad i=1,\dots,n,$ (5.1)
where the latent $\bm{z}_{i}$ are related by some latent causal graph $\mathsf{G}$ and the nonlinearity $f$ is typically estimated with deep neural networks, as in a deep generative model, and causal relationships are inferred through the use of data arising from multiple environments (e.g. across different experimental interventions, time points, etc.). Examples of generative models explicitly of the form (5.1) include diffusion models, variational autoencoders, and flows. Moreover, this class of models continues to dominate leaderboards tracking progress in generative AI (HuggingFace,, 2024; Liu et al.,, 2024).
Remark 1.
The model (5.1) is an assumption on the distribution that can be made independently of whether neural networks are used to fit the data or not; in particular it is not necessary to assume that $f$ itself is a neural network. Of course, in practice we are interested in using neural networks to fit this model nonparametrically.
It is worth contrasting (5.1) with the earlier linear factor model (3.1): The only difference is replacing the linear map $\bm{B}\bm{z}_{i}$ with a nonlinear map $f(\bm{z}_{i})$. Clearly, the former is a special case of the latter. Since even linear factor models are not identifiable, it is clear that (5.1) is also not identifiable. One of the key endeavours in CRL is to determine realizable assumptions under which this model becomes identifiable and estimable.
5.1 Problem definition
To discuss statistical issues such as identifiability, we need to specify the CRL problem a bit more concretely. Given a function class $\mathcal{F}$ and prior family $\mathcal{P}$, we assume the observed data $\bm{x}_{i}\in\mathcal{X}$ is generated from the latent $\bm{z}_{i}\in\mathcal{Z}$ via an unknown mixing function $f\in\mathcal{F}$:
$\bm{x}_{i}=f(\bm{z}_{i})+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}p(\bm{\varepsilon}),\quad i=1,\dots,n,$ (5.2)
where the latent $\bm{z}_{i}\sim p_{\bm{z}}\in\mathcal{P}$ are generated according to a structural equation model:
$\bm{z}_{i}=A\bm{z}_{i}+D^{1/2}\bm{\nu}_{i},\quad\nu_{ik}\perp\!\!\!\perp\operatorname{pa}(z_{ik}),$ (5.3)
where $A$ is the SEM matrix for some DAG $\mathsf{G}$ (cf. (4.2)) and $D$ is a diagonal matrix with positive entries. This prior on $\bm{z}_{i}$ is denoted by $p_{\bm{z}}\in\mathcal{P}$.
Here, $\mathcal{F}$ is a class of transformations and $\mathcal{P}$ is a family of distributions over $\bm{z}$. For example, in the linear factor analysis model $\mathcal{F}$ is a class of linear maps and $\mathcal{P}$ is typically taken to be the family of Gaussian distributions. The model can then be formalized as the collection of distributions induced by these families, and more general models can be considered by simply expanding $\mathcal{F}$ to include nonlinear transformations or $\mathcal{P}$ to include non-Gaussian distributions. A common choice for $\mathcal{F}$ is the family of invertible and differentiable functions; see Kivva et al., (2022) for a detailed discussion of the implications of various functional restrictions.
The goal is to infer the unknown function $f$ and the prior on the latents $p_{\bm{z}}$ including $A$, which encodes the causal graph of the latent variables. The parameters $(f,p_{\bm{z}},A)$ are important for sampling from the model and understanding the effect of interventions on the observed data, with $A$ being important to understand the causal mechanisms and dependence structure.
Remark 2.
In this review, we focus on models where the latent variables satisfy a linear structural equation model (5.3). Of course, one can consider more general models for $p_{\bm{z}}$ (Jiang and Aragam,, 2023; Varici et al.,, 2023, 2024; von Kügelgen et al.,, 2023); see the end of Section 6 for a brief overview.
Remark 3.
The model (5.1) is a nonlinear factor model, which has a long history in both statistics and ML; see Amemiya and Yalcin, (2001) for a review. See also Zhou et al., (2024); Xu et al., (2023) for recent developments. Although the mathematical models share similarities, what differentiates CRL from prior work on nonlinear factor analysis is the focus on deep learning and causality that drives much of CRL. For example, a key distinction arises from the latent SEM (5.3) and the causal, interventionist semantics it introduces.
5.2 Generation vs. understanding
As noted in the introduction, in applications of generative AI there is a crucial distinction between the problems of generation and understanding that parallels the two cultures of statistical modeling. Since the vast majority of the literature on generative AI focuses on generation, we pause for a moment to outline in more detail these two goals:
  1. 1.
    Generation (density estimation, sampling). Here, we evaluate a model based on its ability to reconstruct the true data density of $\bm{x}_{i}$, and ignore what latents are learned or how they can be interpreted. More formally, generation quality is measured by distributional metrics such as total variation, Wasserstein distance, or Kullback-Liebler (KL) divergence. For example, if $p$ represents the density of $\bm{x}_{i}$, then we can evaluate the quality of a model $\widehat{p}$ with the KL divergence $\textup{KL}(p\,\|\,\widehat{p})$. Closely related to this goal is sampling from $p$, which is a key objective in many applications.
  2. 2.
    Understanding (parameter estimation, learning). Here, we evaluate a model based on the quality of the latents themselves, which are captured by the model parameters $\bm{z}_{i}$ and $f$. This could mean recovering the precise latent codes $\bm{z}_{i}$ for each sample $\bm{x}_{i}$, the latent causal graph $\mathsf{G}$, or even just the distribution over $\bm{z}_{i}$. In other words, our model should capture meaningful semantics about the generative process.
The distinction between these two goals has two important statistical implications: 1) Generation is significantly simpler than understanding, and can be achieved without understanding; and 2) To understand models, parameters should be identifiable so that they can be consistently estimated as in traditional statistical models.
5.3 Identifiability
It is clear that (5.1) is not identifiable: Different parameters $(f,p_{\bm{z}})$ lead to the same distribution $p(\bm{x}_{i})$ over the observables. For generation, identifiability is not necessary, however, as we have discussed, we want to go beyond generation to deeper questions of understanding and interpretability. Statistically, we would like to discuss parameter estimation in (5.1), analogous to how factors are estimated in classical linear factor models. While standard in the statistics literature, the relevance of identifiability for ML/AI applications has also been articulated in recent work (D’Amour et al.,, 2020; Reizinger et al.,, 2024).
We shall distinguish between two types of identifiability in the model (5.1):
  1. 1.
    Statistical identifiability of the factor model parameters $(f,p_{\bm{z}})$.
  2. 2.
    Causal identifiability of the causal graph $\mathsf{G}$ and its parameters $A$.
Statistical identifiability can be understood in analogy with factor analysis: This is identifiability of the latent factors themselves. Causal identifiability refers to the distinct problem of identifying the causal relationships between the latent factors as encoded by the latent causal graph $\mathsf{G}$. Of course, both are desirable in practical applications.
These notions of identifiability are distinct: Statistical identifiability is insufficient to imply causal identifiability, and vice versa. That $\mathsf{G}$ cannot be identified from $p_{\bm{z}}$ is a well-known result from graphical models: $\mathsf{G}$ can only be identified up to Markov equivalence. For example, consider two variables $(\bm{z}_{i1},\bm{z}_{i2})$ with joint distribution $p_{\bm{z}}(\bm{z}_{i1},\bm{z}_{i2})$. By the chain rule of probability, $p_{\bm{z}}$ admits two factorizations:
$p_{\bm{z}}(\bm{z}_{i1})p_{\bm{z}}(\bm{z}_{i2}\,|\,\bm{z}_{i1})\quad\text{and}\quad p_{\bm{z}}(\bm{z}_{i2})p_{\bm{z}}(\bm{z}_{i1}\,|\,\bm{z}_{i2}).$
That is, $p_{\bm{z}}$ factorizes with respect to both the graph $\bm{z}_{i1}\to\bm{z}_{i2}$ and $\bm{z}_{i2}\to\bm{z}_{i1}$. Consequently, $\mathsf{G}$ is not identifiable from $p_{\bm{z}}$. The converse that $p_{\bm{z}}$ cannot be identified from $\mathsf{G}$ is a classical result; see Drton et al., (2023) for recent developments.
6 Identifiability in causal representation learning
In this section, we consider in detail the identifiability of the model (5.2-5.3). First, we review approaches to statistical identification of the model parameters $(f,p_{\bm{z}})$. Then, we review approaches to causal identification, i.e. learning the causal graph between the latent variables $Z$ using interventions.
6.1 Statistical identifiability: Nonlinear representation learning
In Section 3, we reviewed the identification problem in linear factor analysis. Of course, identification becomes harder if we relax linearity: Informally, we can always re-write (5.2):
[…]
Thus, without restrictions on $\mathcal{F}$ or $\mathcal{P}$, the model is trivially nonidentifiable. Once we impose restrictions on the parameter spaces $\mathcal{F}$ and $\mathcal{P}$, however, identifiability becomes trickier.
To clarify this, let $m_{\#}p_{\bm{z}}$ denote the push-forward measure of $p_{\bm{z}}$ by $m$, i.e. the distribution on $\bm{z}_{i}$ induced by $m$. Then if (i) $f\circ m^{-1}\in\mathcal{F}$ and (ii) $m_{\#}p_{\bm{z}}\in\mathcal{P}$ for some non-identity $m$, then the model is not identifiable. Following the terminology of Xi and Bloem-Reddy, (2023), we refer to such transformations $m$ that satisfy both (i) and (ii) as indeterminancy transformations; the set of indeterminancy transformations is the indeterminancy set.
Definition 3.
[…]
It is clear that the identity transformation $m=\mathrm{id}$ is always in the indeterminacy set $\mathcal{A}(\mathcal{F},\mathcal{P})$. When $\mathcal{A}(\mathcal{F},\mathcal{P})$ contains additional transformations, then the model parameters are not identifiable. To resolve identifiability issues, we can again further restrict the function class $\mathcal{F}$, the prior distributions $\mathcal{P}$, or both. We state this more precisely in Definition 4; for a formal measure-theoretic treatment of this heuristic, see Xi and Bloem-Reddy, (2023).
Definition 4.
[…]
  1. 1.
    For the model parameterized by $(f,p_{\bm{z}})\in\mathcal{F}\times\mathcal{P}$, we say that $f$ is identifiable up to $\mathcal{A}(\mathcal{F},\mathcal{P})$.
  2. 2.
    In particular, the model parameters $(f,p_{\bm{z}})$ are strongly identifiable if $\mathcal{A}(\mathcal{F},\mathcal{P})=\{\mathrm{id}\}$; that is, the identity function on $\mathcal{Z}$ is the only indeterminancy transformation.
In general, obtaining sufficient conditions for identifying $(f,p_{\bm{z}})$ is easy; we can simply place many restrictions on $\mathcal{F}$ or $\mathcal{P}$. In the linear case, for example, we can constrain the loadings matrix to be upper triangular. The difficulty is finding sufficient conditions that identify the model while remaining as flexible as possible.
In many applications, we may relax the requirement of strong identifiability. For example, we are often not interested in the ordering of the latent variables and so the indeterminancy set can contain permutations of the labels (“label switching”). Other common indeterminacies include scaling (diagonal matrices), translations, and orthogonal transformations.
A common goal in machine learning is “disentanglement” (Bengio,, 2013), which roughly means to separate distinct latent factors of variation. Definition 5 below discusses the relationship between identifiability and “disentangling” latent factors of variation.
Definition 5 (Disentanglement).
[…]
$(\widehat{z}_{\sigma(1)},\dots,\widehat{z}_{\sigma{(K)}})=(g_{1}(z_{1}),\dots,g_{K}(z_{K})),$ (6.2)
where $\sigma$ is a permutation and $g_{k}$ are invertible scalar functions (Ahuja et al.,, 2022; Yao et al., 2024a, ). This is equivalent to the indeterminancy set $\mathcal{A}(\mathcal{F},\mathcal{P})$ containing only element-wise transformations and permutations.
Remark 4.
In the earlier ML literature, “disentanglement” referred to finding independent factors of variation (e.g. Bengio,, 2013). Unfortuantely independence is not sufficient for identification, however, a fact that is well-known in the ICA literature (Hyvärinen and Pajunen,, 1999). This was emphasized in the disentanglement context by Locatello et al., (2019). Consequently, the field moved towards identification-based definitions of disentanglement. Eventually, dependent factors of variation began to be considered, leading to the development of causal representation learning.
6.1.1 Identifiability assumptions on the mixing function
We now review approaches which place assumptions on the mixing function class $\mathcal{F}$ to identify the model.
Sparsity and anchor features.
In the linear case, we saw how the “anchor features” assumption can eliminate rotation indeterminancies. Moran et al., (2022) extend the anchor feature assumption to the nonlinear case by introducing a masking variable $\bm{w}_{j}\in\{0,1\}^{K}$ for each feature $j\in[D]$:
$x_{ij}=f_{\theta}(\bm{w}_{j}\odot\bm{z}_{i})+\varepsilon_{i},$ (6.3)
where $\odot$ denotes the element-wise product. The masking variable $\bm{W}=[\bm{w}_{1}^{T},\cdots,\bm{w}_{D}^{T}]$ is analogous to the sparse loadings matrix in the linear setting. With the anchor assumption, and additional assumptions on the factor covariance and parameters of $f$, Moran et al., (2022) prove that the latent factors are identifiable up to element-wise transformations and permutations. See also Xu et al., (2023), who obtain identifiability with an anchor feature assumption, as well as an additive assumption on $f$ along with independence of latent factors.
Subset condition.
Around the same time, Kivva et al., (2021) proposed the “subset condition” for identifying discrete latent variables, and showed that it is nearly necessary for identification. The subset condition requires that $\operatorname{ch}(\bm{z}_{ij})$ is not a subset of $\operatorname{ch}(\bm{z}_{ik})$ for any $k\neq j$, which can be interpreted as a sparsity condition on $f_{\theta}$. In particular, anchor features are a special case of the subset condition. The basic idea of using sparsity for identifiability dates back to Pearl and Verma, (1992); Evans, (2016).
There have also been recent works on nonlinear ICA which exploit sparsity in $f_{\theta}$ for identifiability (Zheng et al.,, 2022; Zheng and Zhang,, 2023), although ICA models do not consider dependence between the latent factors. These works assume a “structural sparsity” constraint on $f_{\theta}$, which is equivalent to the subset condition of Kivva et al., (2021). Also in the independent factors setting, Markham and Grosse-Wentrup, (2020); Markham et al., (2024) prove identifiability if a certain dependence graph over $\bm{x}_{i}$ admits a unique minimum edge clique cover, for which the anchor feature assumption is sufficient but not necessary.
Independent mechanisms.
The principle of independent mechanisms (Section 4.3) is also related to sparsity-based approaches as well as classical assumptions from PCA. Gresele et al., (2021) operationalize this principle by assuming that the mixing function $f_{\theta}$ has a Jacobian with orthogonal columns. Gresele et al., (2021) prove that IMA eliminates certain classes of mixing functions from the set of indeterminacy transformations, $\mathcal{A}(\mathcal{F},\mathcal{P})$ when $K=D$; Ghosh et al., (2023) extend this result to $K\ll D$. In the linear case, the IMA corresponds to orthogonality conditions on the loadings matrix, as in PCA.
6.1.2 Identifiability assumptions on the latent factors
An alternative strategy to identify the model is to restrict the factor distribution $p_{\bm{z}}$.
Auxiliary information.
A popular approach is to assume there is auxiliary information $\bm{u}_{i}$, which helps to “pin down” $p_{\bm{z}}$. For example, Khemakhem et al., (2020) assume that in addition to $\bm{x}_{i}$, they observe another variable $\bm{u}_{i}$ such that $p(\bm{z}_{i}|\bm{u}_{i})$ lies in an exponential family. This restriction identifies $p_{\bm{z}}$ up to a linear transformation, which can be improved to element-wise transformations when the exponential family has sufficient statistic dimension greater than two.
Mixture priors.
Kivva et al., (2022) eliminate the need for auxiliary data; instead, they assume $p_{\bm{z}}$ is a Gaussian mixture. When the clusters are distinct, this assumption prevent arbitrary rotations of $p_{\bm{z}}$, identifying the factors up to element-wise linear transformations (and permutations). Kivva et al., (2022) also prove identifiability of $f$ up to affine transformations, assuming $f$ is piecewise affine and injective (and $p_{\bm{z}}$ is a Gaussian mixture).
Temporal dependence.
Another approach is to exploit temporal structure in the latent variables. Hälvä et al., (2021) prove that $f$ is identifiable up to element-wise transform and permutation, assuming the latent distributions are stationary. More generally, Ahuja et al., (2021) consider how temporal dependencies restrict the indeterminancy set. Lachapelle et al., (2022) assume sparsity in the latent transition graph to prove element-wise and permutation identifiability. Lippe et al., (2022) assume access to known latent interventions at each time step to prove element-wise and permutation identifiability of $f$; Lippe et al., (2023) allow for unknown binary interventions.
Invariance.
Finally, identifiability can also be achieved by exploiting the existence of invariant structure in the latents under data augmentations or multiple data views. Von Kügelgen et al., (2021) assume that access to augmentations of the observed data $(\bm{x},\widetilde{\bm{x}})$ where a subset of the underlying factors for $\bm{x}$ and $\widetilde{\bm{x}}$ are the same (i.e., fixing the distribution of $\bm{z}$ in some way); this is extended to multiple augmentations in Eastwood et al., (2023). Yao et al., 2024b assume access to multiple views of the data; similarly to data augmentations, the latent variables can also be identified in this case.
6.2 Causal identifiability through interventional data
So far, we have considered the statistical identifiability of $(f,p_{\bm{z}})$. We now move onto causal identifiability of $\mathsf{G}$ and/or $A$, which requires richer types of data such as interventions. As discussed in Section 5.3, we generally cannot infer $\mathsf{G}$ from observational data alone. In fact, even when $\bm{x}_{i}$ is linear in $\bm{z}_{i}$, Squires et al., (2023) show that it is impossible to recover $\mathsf{G}$ without at least one intervention on each latent factor.
This more interesting regime with interventions is called interventional causal representation learning. We now review identifiability results in interventional CRL in (i) the linear mixing case $\bm{x}_{i}=\bm{B}\bm{z}_{i}$; and (ii) the nonlinear mixing case $\bm{x}_{i}=f(\bm{z}_{i})$.
Linear mixing.
Squires et al., (2023) consider the linear mixing case where $\mathcal{F}$ is the set of full rank linear transformations, parametrized by matrices in $\mathbb{R}^{D\times K}$:
$\bm{x}_{i}=\bm{B}\bm{z_{i}}+\bm{\varepsilon}_{i},i=1,\dots,N,$ (6.4)
with $\bm{z}_{i}$ satisfying the linear SEM (5.3) and where $\bm{\varepsilon}_{i}$ has independent components. (Note that $\bm{z}_{i}$ does not need to be Gaussian).
[…]
The key idea behind the identification strategy of Squires et al., (2023) is that the source nodes of the graph $\mathsf{G}$ (i.e. the latent variables without any parents) leave observable implications in the data $\{\bm{x},\bm{x}^{(e)}\}$. Specifically, if $\bm{x}^{(e)}$ corresponds to a source node intervention, then the difference in the precision matrices of $\bm{x}$ and $\bm{x}^{(e)}$ will be rank-one. This is because if a source node is intervened upon, all the downstream dependencies will be the same as the observational data; only the source node distribution is modified. This strategy can be applied iteratively to identify the partial ordering (i.e. topological sort) of the latent causal graph. Now, the precision matrix of $\bm{x}_{i}^{(e)}$ satisfies $\Theta^{(e)}=(B^{\dagger})^{\top}\text{Cov}(z_{i}^{(e)})^{\dagger}B^{\dagger}$, where $\text{Cov}(z_{i}^{(e)})^{\dagger}$ is constrained due to the partial ordering of $\mathsf{G}$. This constraint allows for identifying both $B^{\dagger}$ and $\text{Cov}(z_{i}^{(e)})^{\dagger}$ from $\{\Theta,\Theta^{(e)}\}_{e\in\mathcal{E}}$. [^2] Here, $\dagger$ denotes the Moore-Penrose pseudoinverse.
Nonlinear mixing.
Buchholz et al., (2023) generalize the results of Squires et al., (2023) to the nonlinear setting; specifically, they consider nonlinear mixing functions $f$ which are injective and differentiable. For the latent factors $\bm{z}_{i}$, Buchholz et al., (2023) also consider a linear SEM (5.3), but restrict $\bm{z}_{i}$ to be Gaussian. Buchholz et al., (2023) prove element-wise identifiability of the latent $\bm{z}_{i}$ as well as the causal graph of $\bm{z}_{i}$, again with the assumption of single-node perfect interventions on all the latent variables. The key idea is that the log-odds of $\bm{x}_{i}$ with respect to $\bm{x}_{i}^{(e)}$ is essentially $f^{-1}(\bm{x}_{i})^{T}(\text{Cov}(\bm{z}_{i}^{(e)})-\text{Cov}(\bm{z}_{i}))f^{-1}(\bm{x}_{i})$. Again, the difference between observational and source node precision matrices is rank-one; this geometric structure allows for $f$ to be identified up to linear transformations, at which point the results of Squires et al., (2023) can be applied.
Zhang et al., (2023) also consider nonlinear mixing functions and a linear SEM on $\bm{z}_{i}$, but they allow for soft interventions instead of perfect interventions on $\bm{z}_{i}$. As soft interventions do not remove connections to parent variables, the causal graph can only be identified up to ancestors (more formally, the transitive closure; Tian and Pearl,, 2001). Consequently, Zhang et al., (2023) need an additional sparsity assumption on the graph for identifiability.
Beyond linear SEM.
So far we have reviewed papers with linear SEMs on $\bm{z}_{i}$, although linearity is not strictly needed. Jiang and Aragam, (2023) study a general setting with no parametric assumptions on $\bm{z}_{i}$ and provide a minimal set of nonparametric assumptions under which the latent causal graph $\mathsf{G}$ can be identified up to so-called “isolated edges”. These assumptions are necessary in the sense that if any single assumption is violated, then counterexamples to identifiability can be constructed.
One consequence of avoiding functional constraints is that statistical identifiability becomes challenging. To overcome this, Varici et al., (2024); von Kügelgen et al., (2023) require at least two hard interventions per latent node as well as “interventional discrepancy”, a requirement that the interventional distributions differ almost everywhere. von Kügelgen et al., (2023) require knowing which two interventions operate on the same latent variable; Varici et al., (2024) do not need to know which interventions are paired. The results of Varici et al., (2024); von Kügelgen et al., (2023) rely on the difference in (Stein) score functions of observational and interventional data; for estimated $\widehat{\bm{z}}=h(\bm{x})$, then $\nabla_{\widehat{z}}\log p(\widehat{z})-\nabla_{\widehat{z}}\log p^{(e)}(\widehat{z})=\bm{J}_{h^{-1}}(\widehat{z})[\nabla_{x}\log p(x)-\nabla_{x}\log p^{(e)}(x)]$. With two interventions per latent, Varici et al., (2024) learn $h$ by comparing same-node intervention score differences. Having learned $h$, the latent $\bm{z}_{i}$ and their graph can then be identified.
A different line of work considers an “independent support” condition to identify latent variables (Wang and Jordan,, 2024; Ahuja et al.,, 2023). These papers make an overlap assumption in the latent space; that is, there is non-zero probability on all combinations of latent factors. To provide intuition for this assumption, it is helpful to take an experimental design perspective; we need a design table where all cells (combinations of latent factor levels) have observations. The overlap assumption means that the support of the latent factors is independent, or rectangular. Then, learning a $h$ which has independent support on $\bm{z}$ provably element-wise identifies $\bm{z}$.
7 Practical considerations
In this section, we review current methodological approaches to provide an overview of how CRL can be used in practice.
Variational autoencoders.  A popular approach to fit latent nonlinear representations is the variational autoencoder (Kingma and Welling,, 2014; Rezende et al.,, 2014). Consider the deep generative model (DGM):
$\bm{z}_{i}$ $\sim N(0,I)$
$\bm{x}_{i}$ $=f_{\theta}(\bm{z}_{i})+\varepsilon_{i},\quad\varepsilon_{i}\sim N(0,I).$ (7.1)
A variational autoencoder approximates the latent posterior distribution $p(\bm{z}_{i}|\bm{x}_{i})$ with the variational density:
$q(\bm{z}_{i}|\bm{x}_{i})\sim N(\mu_{\phi}(\bm{x}_{i}),\sigma_{\phi}^{2}(\bm{x}_{i})),$
[…]
As in standard variational inference, the evidence lower bound (ELBO) on the marginal log-likelihood is optimized with respect to the parameters $\{\theta,\phi\}$:
$\log p(\bm{x})\geq\sum_{i=1}^{N}\mathbb{E}_{q_{\phi}(\bm{z}_{i}|\bm{x}_{i})}[\log p_{\theta}(\bm{x}_{i}|\bm{z}_{i})]-D_{\mathrm{KL}}(q_{\phi}(\bm{z}_{i}|\bm{x}_{i})||p(\bm{z}_{i})).$
The first term of the ELBO involves an intractable expectation; this expectation is typically approximated with $L$ Monte Carlo samples of the latent variables:
$\bm{z}_{i}^{(l)}=\mu_{\phi}(\bm{x}_{i})+\sigma_{\phi}(\bm{x}_{i})\odot\varepsilon_{i}^{(l)},\quad\varepsilon_{i}^{(l)}\sim N(0,I),\quad l=1,\dots,L.$
Then, the approximate ELBO is optimized over $\{\theta,\phi\}$ with stochastic gradient ascent.
Identifiable representation learning.  The DGM in (7.1) is not identifiable. As discussed in Section 6.1.1, Moran et al., (2022) introduce a masking variable $\bm{W}$ (6.3); when an anchor assumption holds, the model is identifiable. In practice, a sparsity-inducing prior $p(\bm{W})$ can be placed on the masks. The resultant ELBO is:
$\mathcal{L}(\theta,\phi,\bm{W},\bm{\rho})$ $=\sum_{i=1}^{N}\mathbb{E}_{q_{\phi}(\bm{z}_{i}|\bm{x}_{i})}[\log p_{\theta}(\bm{x}_{i}|\bm{z}_{i},\bm{W})]-D_{\mathrm{KL}}(q_{\phi}(\bm{z}_{i}|\bm{x}_{i})||p(\bm{z}_{i}))+\log p(\bm{W}).$
When the prior $p(\bm{W})$ is hierarchical (e.g. the Spike-and-Slab Lasso of Ročková and George,, 2018), the additional prior parameters may be estimated using the EM algorithm (Moran et al.,, 2022). This sparse VAE can be viewed as a nonlinear analogue to factor analysis; for each column of $\widehat{W}$, the non-zero features be interpreted as a cluster.
Interventional causal representation learning.  Zhang et al., (2023) propose a method for causal representation learning from interventional data, the CausalDiscrepancy VAE. The latent $\bm{z}_{i}$ follow a linear SEM, and the observed data is a nonlinear function of the latent variables:
$\bm{z}_{i}$ $=(1-A)^{-1}\bm{\nu}_{i},\quad\bm{\nu}_{i}\sim N(0,I)$ (7.2)
$\bm{x}_{i}$ $=f_{\theta}(\bm{z}_{i})+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\sim N(0,I),$ (7.3)
where $A\in\mathbb{R}^{K\times K}$ captures the (unknown) causal graph among the $\bm{z}_{i}$. The CausalDiscrepancy VAE estimates the exogenous noise $\bm{\nu}_{i}$ distribution with the variational approximation $q(\bm{\nu}_{i}|\bm{x}_{i})\sim N(\mu_{\phi}(\bm{x}_{i}),\sigma_{\phi}^{2}(\bm{x}_{i}))$, from which $\bm{z}_{i}$ is obtained as $\widehat{\bm{z}}_{i}=(1-\widehat{A})^{-1}\widehat{\bm{\nu}}_{i}$.
[…]
$\mathcal{L}(\theta,\phi,\varphi,A)$ $=\sum_{i=1}^{N}\mathbb{E}_{q_{\phi}(\bm{\nu}_{i}|\bm{x}_{i})}[\log p_{\theta}(\bm{x}_{i}|\bm{z}_{i})]-D_{\mathrm{KL}}(q_{\phi}(\bm{\nu}_{i}|\bm{x}_{i})||p(\bm{\nu}_{i}))$ (7.4)
$\quad-\mathrm{MMD}(P_{\widehat{\theta},\widehat{\varphi}}(\widehat{X}^{(e)})||X^{(e)})-\lambda\lVert A\rVert_{1}.$ (7.5)
Zhang et al., (2023) apply the CausalDiscrepancy VAE to Peturb-Seq data (Norman et al.,, 2019). In addition to an observational dataset, there are interventional datasets, each corresponding to a single-gene CRISPR activation. The function $T_{\varphi}$ maps the intervention label (gene label) to a latent dimension; the latent dimensions can then be interpreted by inspecting which gene interventions are mapped to the same latent variable.
Score-based methods.  In Section 6.2, we discussed papers which use score functions of data to identify latent causal graphs. Two popular approaches for estimating score functions are (i) using Stein’s identity (Stein,, 1972; Rolland et al.,, 2022), and (ii) sliced score matching with variance reduction (Song et al.,, 2020).
Given the estimated data score functions, Varici et al., (2024) note that
$\nabla_{\widehat{z}}\log p(\widehat{z})-\nabla_{\widehat{z}}\log p^{(e)}(\widehat{z})=\bm{J}_{h^{-1}}(\widehat{z})[\nabla_{x}\log p(x)-\nabla_{x}\log p^{(e)}(x)],$ (7.6)
for any function $h$ with $\widehat{z}=h(X)$. These differences between the scores of the interventional and observational data should be small because the interventions act on a single-node at a time. Consequently, Varici et al., (2024) learn an $h$ that minimizes the differences in score functions. This $h$ is then used to obtain the latent estimates; the latent causal graph can also be obtained from the score function differences.
8 Conclusion and open questions
We have reviewed CRL and its connections to classical statistical methods such as factor analysis and structural equation modeling. CRL is an emerging discipline with roots in machine learning, statistics, and causality, and draws inspiration from a variety of domains to address important practical challenges in the use of generative models in scientific applications where interpretability and causality are key. From a statistical perspective, CRL offers a new twist on classical factor analysis through its principled use of causal semantics (graphs, interventions, etc.).
Although the ideas behind CRL have evolved slowly over many decades, it is still a developing discipline with numerous open questions. We conclude this review with a brief summary of some directions for future inquiry.
Beyond interventions
In this review, we have focused on using interventions for identification. While this is indeed practical in a variety of settings, there is significant interest in generalizing beyond interventions. An emerging paradigm is to use multiple environments (i.e. non-iid data arising from data augmentation, distribution shifts, different locations, times, contexts, etc.) to identify latent structure and/or improve generalization (Peters et al.,, 2016; Huang et al.,, 2020; Perry et al.,, 2022; Mooij et al.,, 2020). Indeed, experimental interventions are a special class of environments. It is an intriguing problem to understand how general environments can enable identifiability when strict interventions are not available; see e.g. Varici et al., (2024); Chen et al., (2024).
Implementation and evaluation
Early work on CRL has mostly focused on the identification problem, however, just as important is the problem of estimation and practical algorithms. Despite substantial advances in practical aspects of training generative models, these advances have not directly translated over to the setting of CRL. Recent work (Section 7) has begun to address these challenges, but no general framework has emerged. Another important problem is evaluating CRL models: Since we can only identify parameters up to indeterminacy transformations, how can we assess the quality of estimators in a principled way? One approach is to use downstream tasks such as out-of-distribution generalization (Lu et al.,, 2021) and extrapolation (Saengkyongam et al.,, 2024).
Statistical theory
There are significant gaps in our statistical understanding of CRL: Finite-sample rates, misspecification, uncertainty quantification, and model selection all remain underexplored. Some recent exceptions in this direction include Acartürk et al., (2024); Buchholz and Schölkopf, (2024), although many problems remain: Since the models are nonparametric, the usual curse of dimensionality applies, but can causal assumptions help mitigate the curse of dimensionality? Alternatively, if we aim only to estimate the graph $\mathsf{G}$ and its parameters $A$, are conventional rates achievable? How can traditional estimation strategies such as variational inference lead to consistent model selection, and what are its tradeoffs? What about robustness and estimation under misspecification? These are just some of the intriguing statistical questions that demand resolution.
Large language models
The development of large language models (LLMs) and more generally, foundation models, has occurred in parallel with CRL. This suggests the tantalizing possibility of developing pre-trained foundation models with causal interpretability. This comes with numerous practical challenges; e.g. connecting the objectives of next-token prediction with causal objectives; principled use of diverse, multi-modal, internet-scale datasets; extracting interpretable concepts; etc. On the latter, recent work has made progress (Rajendran et al.,, 2024). An alternative line of work extracts interpretable concepts post-hoc from pre-trained LLMs using sparse dictionary learning (Bricken et al.,, 2023; Templeton et al.,, 2024). These sparse dictionary learning approaches assume that each token is the linear combination of a small set of concept vectors, selected from a overcomplete concept dictionary. That is, the concept space has a much higher dimension than the data, in contrast to most CRL methods. Investigating these higher-dimensional latent space settings in the context of CRL is an important area for future research.